# The Detection of KIC 1718360, A Rotating Variable with a Possible Companion, Using Machine Learning

### v2

@@ Introduction
This paper presents the finding of a periodic dip in the lightcurve of the star KIC 1718360 that was detected in both Kepler and TESS data using a One-Class Support Vector Machine (SVM) model. This observation is indicative of rotational variability in the G1.5IV-V type star. The star is thought to have a size comparable to that of the sun, with an estimated size of 0.91 $R_{\odot}$
@@
The parameters in Table 2 show that this exoplanet candidate, if it exists, must orbit in extreme proximity to its host star. Additionally, it seems to be extremely small, approximately the size of Mercury. However, this is not without precedent1010footnotemark: 101111footnotemark: 11.
@@ Conclusions
Additionally, it seems there is a case to be made for the existence of a possible exoplanetary companion orbiting the star. In addition to the main dip caused by rotational variability, there seems to be a periodic secondary dip in the lightcurve of KIC 1718360, which could suggest the existence of a Mercury-sized exoplanet orbiting this star. Further study is needed to confirm or deny the existence of this exoplanet candidate.

### v3

@@ Introduction
This paper presents the finding of a periodic dip in the lightcurve of the star KIC 1718360 that was detected in both Kepler and TESS data using a One-Class Support Vector Machine (SVM) model. This observation is indicative of rotational variability in the G1.5IV-V type star. The star is thought to have a size comparable to that of the sun, with an estimated size of 1.09211 $R_{\odot}$
@@
The parameters in Table 2 show that this exoplanet candidate, if it exists, must orbit in extreme proximity to its host star. However, this is not without precedent1010footnotemark: 10. Additionally, it seems to be in the expected radius range of a super-Earth.
@@ Conclusions
Additionally, it seems there is a case to be made for the existence of a possible exoplanetary companion orbiting the star. In addition to the main dip caused by rotational variability, there seems to be a periodic secondary dip in the lightcurve of KIC 1718360, which could suggest the existence of a super-Earth orbiting this star. Further study is needed to confirm or deny the existence of this exoplanet candidate.